please help me! can someone pls? :(
solve the equation. check your answers.
3/m-2 + 1/m-1 = 7/m^2-3m+2

Answer by stanbon(75887)

3/m-2 + 1/m-1 = 7/m^2-3m+2
----
Multiply thru by m^2-3m+2 which equals (m-2)(m-1)
-----
= 3(m-1) + (m-2) = 7
----
= 3m-3+m-2 = 7
-----
4m -5 = 7
4m = 12
m = 3
==============
Checking:
3/m-2 + 1/m-1 = 7/m^2-3m+2
3/(3-2) + 1/(3-1) = 7/(3^2-3*3+2)
---
3/1 + 1/2 = 7/2
----
3 1/2 = 3 1/2
==================
Cheers,
stan H.
==================